# Proposal for a 3-Wave Mixing Element with Quantum Paraelectric Materials

### v1

@@ Abstract
At cryogenic temperatures and microwave frequencies, the perovskite crystals strontium titanate (STO) and potassium tantalate (KTO) have large, tunable permittivity arising from a quantum paraelectric phase. As such, these materials hold promise as a platform to realize compact, variable capacitance elements for use in quantum devices. From modulating this capacitance, we propose the development of a parametric mixing element: a quantum “paraelectric nonlinear dielectric amplifier” (PANDA). We calculate that a PANDA made from a nanofabricated parallel plate capacitor and realistic design constraints can demonstrate a three-wave mixing strength of order MHz, in comparison to an effective Kerr strength of sub-Hz. This suggests excellent performance as a three-wave mixing element, with high compression power in analogy to superconducting parametric amplifiers based on kinetic inductance. Beyond parametric amplifiers, we predict that compact, tunable capacitors based on STO, KTO, and related materials can enable a wide class of cryogenic quantum circuits including novel filters, switches, circulators, and qubits.
@@
The growth of quantum technology has created the need for cryogenic circuit elements that are compact, tunable, and low-loss. Superconducting quantum processors, for example, are a leading platform for quantum information, but operate under a stringent set of requirements such as millikelvin temperatures, limited external fields, and signals at the level of single microwave photons [1]. This motivates the development of novel devices engineered to have improved performance and scalability in these regimes.
@@
Superconducting qubits are based on the Josephson junction (JJ), an inductive nonlinearity. JJs are also the building block for much of the circuitry adjacent to the qubits, including the quantum-limited parametric amplifiers. These amplifiers are required for qubit readout [2, 3], and, have utility for fundamental science, e.g. axion dark matter detection [4]. As such, there is a growing body of literature on improving the gain, bandwidth, power handling, and scalability of parametric amplifiers, all while pushing noise performance as close as possible to the quantum limit [5, 6, 7, 8, 9, 10, 11, 12].
@@
However, JJ-based parametric amplifiers are not suitable for all applications. The compression power, marking the onset of nonlinear saturation, is typically limited to below approximately -100 dBm [3]. JJ’s are also susceptible to the breakdowns in superconductivity that occur under high magnetic fields and optical illumination. Yet, there are many applications for quantum-limited measurement under these conditions. Microwave-to-optical transducers require close interplay between microwave circuits and optical fields [13, 14], and magnetic fields are essential to the study of solid-state spin defects [15], axion dark matter haloscopes [16, 17], and Majorana qubits [18], for example.
@@
The rich application space for quantum-limited parametric amplifiers motivates the development of novel parametric mixing elements [19, 6]. Alternative nonlinearities to the Josephson junction include kinetic inductance [20, 21, 22, 23, 24, 25, 26], quantum dots [27], field-effect transistors [28], and hybrid superconductor-semiconductor junctions [29]. In particular, kinetic inductance parametric amplifiers (KIPAs) have recently been shown to achieve high gain with low noise at compression powers of approximately -50 dBm and Tesla-scale magnetic fields [21, 30, 24].
@@
Here, we propose the development of quantum paraelectric nonlinear dielectric amplifiers (PANDAs) as an alternative parametric mixing element for use in quantum technologies. Building on Ref. [31], we analyze the design of a PANDA realized from a tunable capacitor (varactor) made from the perovskites strontium titanate, $\mathrm{SrTiO_{3}}$ (STO) or potassium tantalate, $\mathrm{KTaO_{3}}$ (KTO), quantum paraelectric materials which have voltage tunable dielectric permittivity. We derive the nonlinear Hamiltonian for a resonator made from an STO or KTO capacitor, and calculate that a nanofabricated capacitor geometry can have three-wave mixing strength as great as order 10 MHz with order $\sim 0.1\,\mathrm{Hz}$ effective Kerr terms (a ratio of $10^{8}$). This is far greater nonlinearity than varactors made from bulk substrates [32], and suggests high parametric gain and competitive dynamic range compared to state-of-the-art KIPAs [21, 30, 24]. Overall, our proposal examines the rich design space of RF/microwave circuits made from perovskite dielectrics and related materials, motivates open challenges in the nanofabrication of these materials, and serves as a blueprint for future experiments that seek to realize these emerging cryogenic components.
@@
Dielectric Properties — STO and KTO have a similar structure in which a Ti/Ta atom is centered within an octahedral cage of O atoms, Fig. 1a. In the paraelectric phase, the material becomes polarized when the central atom is displaced by an electric field $E$, changing the dielectric properties. The resulting normalized bias is parameterized by $\lambda=\sqrt{\lambda_{s}^{2}+(E/E_{N})^{2}}\approx E/E_{N}$, where $E_{N}$ is the re-normalizing field and $\lambda_{s}$ is a dimensionless measure of material inhomogeneity, that goes to zero for an ideal crystal. Permittivity and loss depend on bias as described by a well-known modified Landau-Ginzburg-Devonshire (LGD) theory [33, 34, 35, 36, 37], with permittivity approximated by:
@@
Because here we are interested in RF and microwave devices engineered to operate at cryogenic temperatures for quantum applications, we consider the LGD theory in the limit of low frequency (order 100 GHz and below [38, 39]) and low temperature, only. From measurement of bulk crystals at 4 Kelvin and at zero field ($\lambda=0$) these models [36, 37, 40] predict a relative permittivity and loss tangent of $\varepsilon_{r}=\varepsilon/\varepsilon_{0}\approx 24\times 10^{3}$ and $\tan(\delta)\approx 10^{-3}$ for STO, and $\varepsilon_{r}\approx 4.5\times 10^{3}$ and $\tan(\delta)\approx 10^{-5}$ for KTO. In the limit of large field ($\lambda\gg 0$), $\varepsilon_{r}$ decreases to order one thousand, only, while loss increases (Appendix A) Both permittivity and loss can depend on material properties including defect concentration and strain. Minimizing loss in both bulk and thin-films is an open research challenge [41, 42, 43].
Figure 1: (a) The perovskite crystals strontium titanate (STO) and potassium tantalate (KTO) have a cage-like structure where a central Ti or Ta atom lies within a cell formed by O and Sr or K atoms, respectively. At cryogenic temperatures these materials have high permittivity and low loss. Permittivity is tunable via an applied electric field $\vec{E}$, which displaces the central atom and polarizes the crystal. (b) We consider a parallel plate capacitor, with plate area $A$ and separation $d$, and with the dielectric made from a thin film of STO or KTO. A voltage $v$ across the plates induces an electric field $E=v/d$, which changes $\varepsilon(v)$ and $C(v)=\varepsilon_{0}\varepsilon_{r}(v)A/d$. Capacitance is plotted choosing $A=(4\,\mathrm{\mu m})^{2}$ and $d=200\,\mathrm{nm}$, and the dielectric properties given in Appendix A.
@@
A Nonlinear Resonator — We propose to make use of these dielectric properties to develop compact and tunable capacitors made from STO, KTO, or related materials, and which can be integrated into resonant circuits. In this work, we consider a lumped element resonator implemented with a parallel plate capacitor with area $A$ and a STO/KTO dielectric thin film of thickness $d$ with superconducting metallization on the faces (Fig. 1b). However our analysis is also applicable to a coplanar waveguide (CPW) design on a bulk STO/KTO substrate [40], which differs only in the geometric design space. Moreover, we envision a galvanic connection to one of the metal faces, which enables electric field tuning of the material with an external bias voltage.
@@
Figure 2: (a) A resonator is formed by combining an inductor ($L$) and capacitor ($C$). We model that a DC bias voltage $v_{0}$ can be applied across the capacitor plates, changing the capacitance as described in Fig. 1. (b) This nonlinear resonant circuit is analogous to an oscillating mass on a spring whose spring constant stiffens when stretched. We are interested in its dynamics, Eq. 2, given both a static restoring force set by $v_{0}$, and, a parametric driving force set by $v_{\mathrm{ac}}$. (c) The resonant frequency $\omega_{0}/2\pi$ choosing $L=0.5\penalty 10000\ \,\mathrm{nH}$, (d) the three-wave mixing (3WM) strength $\xi/2\pi$ choosing $v_{\mathrm{ac}}=1\,\mathrm{mV}$ (AC bias charge $q_{\mathrm{ac}}=v_{\mathrm{ac}}C(v)$), (e) and the effective Kerr strength $K_{\mathrm{eff}}/2\pi$, all as functions of $v_{0}$.
[…]
@@
A Degenerate Parametric Amplifier— To use our circuit as a degenerate parametric amplifier, we will modulate the charge on the capacitor near twice the resonant frequency $\omega_{p}\sim 2\omega_{0}$ in the presence of an offset charge $q_{0}$ induced by the external bias. We can model the resulting dynamics of small perturbations around the charge equilibrium set by $v_{0}$ via:
@@
with a 3WM strength $\xi=-U_{c}^{(3)}q_{\mathrm{ac}}q_{\mathrm{zpf}}^{2}e^{-i\theta}/2\hbar$. This term is proportional to both the third derivative of the capacitor’s energy vs. charge, $U_{c}^{(3)}(q)=\partial^{3}U_{c}/\partial q^{3}$ where $U_{c}(q)=\int_{0}^{q}v(q^{\prime})dq^{\prime}$, and, to the amplitude of a drive $q_{\mathrm{drive}}(t)=q_{\mathrm{ac}}\cos(2\omega_{0}t+\theta)$ at twice the circuit’s resonance frequency. The effective Kerr term $K_{\mathrm{eff}}=U_{c}^{(4)}q_{\mathrm{zpf}}^{4}/2\hbar$ is proportional to the fourth derivative $U_{c}^{(4)}(q)=\partial^{4}U_{c}/\partial q^{4}$ of energy with respect to charge. In both terms, $q_{\mathrm{zpf}}$ is the capacitor’s zero-point charge fluctuation. As is standard in the study of parametric amplifiers [44, 8, 21], Eq. 2, is derived (Appendix C) within the rotating wave approximation, and, at small enough amplitudes so that higher-order dynamics can be neglected.
@@
The calculations shown in Fig. 2 indicate that a properly engineered and operated PANDA can be an excellent three-wave mixing element with minimal unwanted four-wave mixing (Kerr) terms that degrade performance. Using the capacitor parameters given in Fig. 1b and, as an illustrative example, an AC bias of $v_{\mathrm{ac}}=1\penalty 10000\ \,\mathrm{mV}$, we find a maximum three-wave mixing (3WM) strength of approximately 26 MHz for STO and 9.5 MHz for KTO. Our choice of $v_{\mathrm{ac}}$ translates to pumping with approximately $10^{4}$ photons for both materials; a precise comparison will depend on broader design considerations.
@@
The optimal 3WM points occur at a DC bias of 9.3 mV for STO and 66 mV for KTO, respectively; modest voltages compatible with superconducting electronics. At these operation points $K_{\mathbf{eff}}$ is of order 0.1 Hz for both materials. For comparison, the Josephson parametric amplifier (JPA) in Ref. [9], operated as a linear amplifier, has $\xi/2\pi\lesssim 25\,\mathrm{MHz}$ and $K_{\mathbf{eff}}/2\pi\approx 50\,\mathrm{kHz}$, and the KIPA in Ref. [21] operates at $\xi/2\pi\lesssim 27\,\mathrm{MHz}$ and $K_{\mathbf{eff}}\approx 0.1\,\mathrm{Hz}$. Our maximum predicted figure-of-merit ratio of $\xi/K_{\mathrm{eff}}\approx 10^{8}$ is much higher than the $\xi/K_{\mathrm{eff}}\lesssim 10^{5}$ ratio typical of Josephson parametric amplifiers (JPA) [45, 6, 8] and is comparable to state-of-the-art KIPAs [20, 21, 30, 24].
@@
Figure 3: (a) Input-output model for a parametric amplifier made from an STO or KTO capacitor. Modulation of the capacitance around $\omega_{p}\approx 2\omega_{0}$ yields degenerate parametric amplification. (b,c) Reflection $R(\omega)$ of a STO and KTO device, respectively, plotted for different values of $\xi$ normalized to resonator loss $\kappa/2$. The model (Eq. 3) uses the parameters as given in Fig. 2 and an external quality factor of $Q_{\mathrm{ext}}=100$.
@@
To analyze our circuit as a degenerate parametric amplifier, we consider the reflection of a weak signal tone while the capacitor is charge pumped as previously discussed. For simplicity, we consider a single-port implementation of our amplifier (Fig. 3a). Following an input-output formalism applied to $H_{\mathrm{driven}}$ (Eq. 2), the circuit reflection coefficient is [46, 21],
@@
Here $\Delta=\omega_{0}-\omega_{p}/2$ is the detuning of the half-pump frequency from the resonator, and for remaining analysis we choose $\Delta=0$.
@@
A critical design parameter is the external coupling rate $\kappa_{\mathrm{ext}}$, which should be much larger than the internal loss rate $\kappa_{\mathrm{int}}$ for optimal noise performance. $\kappa_{\mathrm{ext}}$ will be determined by the details of the coupler, whereas $\kappa_{\mathrm{int}}$ depends on the material of choice. Our requirement of introducing both DC and AC bias fields across the capacitor translates to having an inductive coupler. For example, $\kappa_{\mathrm{ext}}$ can be controlled by implementing the coupler as a stepped-impedance filter (see Fig. 6), in further analogy to kinetic inductance amplifiers. For the estimates in this paper, we consider an external quality factor $Q_{\mathrm{ext}}=\omega_{0}/\kappa_{\mathrm{ext}}=100$ for both designs, as is typical for superconducting parametric amplifiers [8, 21, 30, 24].
@@
We expect internal loss to be dominated by bulk dielectric loss. The internal quality factor $Q_{\mathrm{int}}=\omega_{0}/\kappa_{\mathrm{int}}=\tan^{-1}{(\delta)}$ is evaluated at $v_{0}$ to be $Q_{\mathrm{int}}\approx 6.1\times 10^{2}$ for STO and $Q_{\mathrm{int}}\approx 7.4\times 10^{3}$ for KTO (see Fig. 4c). These parameters yield $\kappa_{\mathrm{int}}/2\pi=3.4\,\mathrm{MHz}$ and $\kappa_{\mathrm{ext}}/2\pi=20.7\,\mathrm{MHz}$ for the STO design, and $\kappa_{\mathrm{int}}/2\pi=0.7\,\mathrm{MHz}$ and $\kappa_{\mathrm{ext}}/2\pi=48.9\,\mathrm{MHz}$ for the KTO design. Both designs are over-coupled ($\kappa_{\mathrm{ext}}/\kappa_{\mathrm{in}}>1$), but the KTO design is more overcoupled due to its lower modeled internal loss. This suggests that PANDAs made from KTO may have better noise performance.

### v2

# Three-Wave Mixing Element with Quantum Paraelectric Materials

@@ Abstract
At cryogenic temperatures and microwave frequencies, the perovskite crystals strontium titanate (STO) and potassium tantalate (KTO) have large, tunable permittivity arising from a quantum paraelectric phase. As such, these materials hold promise as a platform to realize compact, variable capacitance elements for use as three-wave mixing elements in quantum devices. From modulating this capacitance within a resonant circuit, we propose the development of a quantum “paraelectric nonlinear dielectric amplifier” (PANDA). We calculate that a PANDA made from a nanofabricated parallel plate capacitor and realistic design constraints can demonstrate a three-wave mixing strength of order MHz, in comparison to an effective Kerr strength of sub-Hz. This suggests excellent performance as a degenerate parametric amplifier, with high compression power in analogy to superconducting parametric amplifiers based on kinetic inductance. Beyond parametric amplifiers, we predict that compact, tunable capacitors based on STO, KTO, and related materials can enable a wide class of cryogenic quantum circuits including novel filters, switches, circulators, and qubits.
I Introduction
@@
The growth of quantum technology has created the need for cryogenic circuit elements that are compact, tunable, and low loss. Superconducting quantum processors, for example, are a leading platform for quantum information, but operate under a stringent set of requirements such as millikelvin temperatures, limited external fields, and signals at the level of single microwave photons [1]. This motivates the development of devices engineered to have improved performance and scalability in these regimes.
@@
Superconducting qubits are based on the Josephson junction (JJ), an inductive nonlinearity. JJs are also the building block for much of the circuitry adjacent to the qubits, including the quantum-limited parametric amplifiers. These amplifiers are required for qubit readout [2, 3], and, have utility for fundamental science, e.g. axion dark-matter detection [4]. As such, there is a growing body of literature on improving the gain, bandwidth, power handling, and scalability of parametric amplifiers, all while pushing noise performance as close as possible to the quantum limit [5, 6, 7, 8, 9, 10, 11, 12].
@@
However, JJ-based parametric amplifiers are not suitable for all applications. The compression power, marking the onset of nonlinear saturation, is typically limited to below approximately -100 dBm [3]. JJs are also susceptible to the breakdowns in superconductivity that occur under high magnetic fields and optical illumination. Yet, there are many applications for quantum-limited measurement under these conditions. Microwave-to-optical transducers require close interplay between microwave circuits and optical fields [13, 14], and magnetic fields are essential to the study of solid-state spin defects [15], axion dark-matter haloscopes [16, 17], and Majorana qubits [18], for example.
@@
The rich application space for quantum-limited parametric amplifiers motivates the development of alternative parametric mixing elements [19, 6]. Alternative nonlinearities to the Josephson junction include kinetic inductance [20, 21, 22, 23, 24, 25, 26], quantum dots [27], field-effect transistors [28], and hybrid superconductor-semiconductor junctions [29]. In particular, kinetic inductance parametric amplifiers (KIPAs) have recently been shown to achieve high gain with low noise at compression powers of approximately -50 dBm and tesla-scale magnetic fields [21, 30, 24].
Parametrically tunable capacitors (varactors) are a conceivable alternative to these nonlinearities. Varactors have long appeared within classical, room-temperature electrical engineering literature [31, 32], including work from the mid-$\mathrm{20^{th}}$ century that makes use of strontium titanate $\mathrm{SrTiO_{3}}$ (STO) [33, 34]. Yet despite this storied history, varactors have thus far received limited attention within the context of quantum information experiments that operate at microwave frequencies and millikelvin temperatures, or that utilize quantum paraelectric phases.
@@
Here, we propose using the tunable capacitance of quantum paraelectric materials as an alternative parametric mixing element for quantum circuits. In particular, we propose the development of quantum paraelectric nonlinear dielectric amplifiers (PANDAs) as an alternative parametric amplifier for use in quantum technologies. Building on Ref. [35], we analyze the design of a PANDA realized from a tunable capacitor (varactor) made from the perovskites strontium titanate, $\mathrm{SrTiO_{3}}$ (STO) or potassium tantalate, $\mathrm{KTaO_{3}}$ (KTO), quantum paraelectric materials which have large voltage tunable dielectric permittivities. We derive the nonlinear Hamiltonian for a resonator made from an STO or KTO capacitor, and calculate that a nanofabricated capacitor geometry can display three-wave-mixing strengths (i.e. parametric coupling rate) as great as approximately $O(10\,\textrm{MHz})$ with approximately $O(0.1)\,\mathrm{Hz}$ effective Kerr terms (a ratio of $10^{8}$). This is greater nonlinearity than varactors made from bulk substrates [36], and suggests high parametric gain and competitive dynamic range compared to state-of-the-art KIPAs [21, 30, 24]. Overall, our proposal examines the rich design space of rf/microwave circuits made from perovskite dielectrics and related materials, motivates open challenges in the nanofabrication of these materials, and serves as a blueprint for future experiments that seek to realize these emerging cryogenic components.
Figure 1: (a) The perovskite crystals strontium titanate (STO) and potassium tantalate (KTO) have a cagelike structure where a central Ti or Ta atom lies within a cell formed by O and Sr or K atoms, respectively. At cryogenic temperatures these materials have high permittivity and low loss. Permittivity is tunable via an applied electric field $\vec{E}$, which displaces the central atom and polarizes the crystal. (b) A voltage tunable resonator is formed by a parallel plate capacitor made from thin-film STO or thin-film KTO, and, a coupling inductance $L$. An external voltage bias $v_{0}$ across the plates induces an electric field $E=v_{0}/d$ and tunes the capacitance. (c) Capacitance and resonance frequency $\omega_{0}=1/\sqrt{LC(v)}$ as a function of $v_{0}$ for $A=(4\,\mathrm{\mu m})^{2}$ and $d=200\,\mathrm{nm}$ with the dielectric properties given in Appendix A. Parameters are chosen to enable microwave frequency operation and a compact footprint.
II Dielectric Properties
@@
STO and KTO have a similar structure in which a Ti/Ta atom is centered within an octahedral cage of O atoms, Fig. 1a. In the paraelectric phase, the material becomes polarized when the central atom is displaced by an electric field $E$, changing the dielectric properties. The resulting normalized bias is parameterized by $\lambda=\sqrt{\lambda_{s}^{2}+(E/E_{N})^{2}}\approx E/E_{N}$, where $E_{N}$ is the re-normalizing field and $\lambda_{s}$ is a dimensionless measure of material inhomogeneity, that goes to zero for an ideal crystal. Permittivity and loss depend on bias as described by a well-known modified Landau-Ginzburg-Devonshire (LGD) theory [37, 38, 39, 40, 41], with permittivity approximated by:
@@
Because here we are interested in rf and microwave devices engineered to operate at cryogenic temperatures for quantum applications, we consider the LGD theory in the limit of low frequency (order 100 GHz and below [42, 43]) and low temperature, only. From measurement of bulk crystals at 4 kelvin and at zero field ($\lambda=0$) these models [40, 41, 44] predict a relative permittivity and loss tangent of $\varepsilon_{r}=\varepsilon/\varepsilon_{0}\approx 24\times 10^{3}$ and $\tan(\delta)\approx 10^{-3}$ for STO, and $\varepsilon_{r}\approx 4.5\times 10^{3}$ and $\tan(\delta)\approx 10^{-5}$ for KTO. In the limit of large field ($\lambda\gg 0$), $\varepsilon_{r}$ decreases to order one thousand and modifies the loss(Appendix A). In reality, both permittivity and loss can depend on material properties including defect concentration and strain. Minimizing loss in both bulk and thin films is an open research challenge [45, 46, 47]. For example, Ref. [48] reports permittivity of approximately $2\times 10^{3}$ and a loss tangent of $3.4\times 10^{-4}$ in thin-film STO at microwave frequencies.
III A Nonlinear Resonator
@@
We propose to make use of these dielectric properties to develop compact and tunable capacitors made from STO, KTO, or related materials, and which can be integrated into resonant circuits. In this work, we consider a lumped-element resonator implemented with a parallel-plate capacitor (PPC) with area $A$ and an STO or KTO dielectric thin film of thickness $d$ with superconducting metallization on the faces (Fig. 1b). However our analysis is also applicable to a coplanar waveguide (CPW) design on a bulk STO or KTO substrate [44], which differs only in the geometric design space. Moreover, we envision a galvanic connection to one of the metal faces, which enables electric-field tuning of the material with an external bias voltage.
@@
Figure 2: (a) Our nonlinear resonant circuit is analogous to an oscillating mass on a spring whose spring constant stiffens when stretched. We are interested in its dynamics, Eq. 2, given both a static restoring force set by $v_{0}$, and, a parametric driving force set by $v_{\mathrm{ac}}$. (b) By Taylor expanding the potential around a variable bias $v_{0}$, we determine the nonlinear coefficients $g_{3}$ and $g_{4}$.
With an eye towards quantum-limited parametric amplification at approximately mK temperatures, we begin by highlighting our circuit design criteria. We target resonance frequencies in the approximately GHz range using the STO or KTO capacitance $C(v)=\varepsilon_{0}\varepsilon_{r}(v)A/d$ (Fig. 1c). This capacitance must be widely tunable to enable commensurate nonlinear mixing capabilities. We neglect any stray capacitance to ground given that $\varepsilon_{r}\gg\varepsilon_{\mathrm{substrate}}$ for conventional substrate hosts (i.e. silicon or sapphire). To tune $C(v)$, it is convenient to work with volt-scale biases compatible with standard low-noise control electronics, which in turn influences the choice of dielectric thickness $d$. As such, varactors made from bulk material are not optimal for our applications [36]. We also limit film thicknesses to >>100 nm, which have been shown to display bulklike low-temperature dielectric properties [49]; thinner films may suffer from dead-layer effects. Finally, design parameters should be compatible with a reasonable fabrication process.
Given these considerations, we choose as an illustrative example an inductance of $L$ = 0.5 nH, and capacitor dimensions of $A=(4\,\mathrm{\mu m})^{2}$ and $d=200\,\mathrm{nm}$ (a released film thickness of 105 nm was recently experimentally demonstrated for STO [46]). This corresponds to a tunable capacitance $C(v)=\partial q/\partial v=A\varepsilon_{0}\varepsilon_{r}(v)/d$ between $17.2$ and $1.28\,\mathrm{pF}$ for STO, and between $3.18$ and $0.86\,\mathrm{pF}$ for KTO using a bias range $\{0,250\}$ mV and the dielectric model in Appendix A. The magnitude of this tunability is comparable to that of inductive elements based on superconducting quantum interference devices (SQUIDs).
The resulting tunable, nonlinear resonator can be described by the following Hamiltonian:
$H/\hbar=\omega_{0}a^{\dagger}a+ig_{3}\big(a-a^{\dagger}\big)^{3}+g_{4}\big(a-a^{\dagger}\big)^{4}+$..., (2)
where $\omega_{0}$ is the resonator frequency, $g_{3}$ and $g_{4}$ are third- and fourth- order nonlinearities of interest, and $a$ ($a^{\dagger}$) are complex variables that can be cast in terms of the bosonic annihilation (creation) operators in the quantum mechanical description (see Appendix C). This Hamiltonian has strong analogy to the inductive mixing elements conventionally used in circuit quantum electrodynamics (QED), e.g. the SNAIL-based circuits in Ref. [6]. The sign of the nonlinear terms in Eq. 2 is based on the standard definition of the quadrature operators [50], and leads to a difference from conventional circuit QED devices because in this case the nonlinearity is capacitive rather than inductive. The parameters $\{\omega_{0},g_{3},g_{4}\}$ all depend on the PPC capacitance, and therefore can be tuned with $v_{0}$. As shown in Fig. 2b, our proposed design can simultaneously achieve a high ratio between the third- and fourth-order nonlinearities, $g_{3}/2\pi\sim O(\textrm{kHz})$ and $g_{4}/2\pi\sim O(\textrm{mHz})$, and holds promise as a compact and high-performance three-wave mixing element for microwave frequency quantum circuits.
IV A Degenerate Parametric Amplifier
@@
Parametric amplification is a key potential use case for this nonlinear resonator (the PANDA). To use our circuit as a degenerate parametric amplifier, we will modulate the charge on the capacitor near twice the resonant frequency $\omega_{p}\sim 2\omega_{0}$ and with phase $\theta$ in the presence of an offset bias $v_{0}$. For simplicity, we consider a single-port implementation of our amplifier (Fig. 3a). We can model the resulting dynamics of small perturbations around the charge equilibrium set by $v_{0}$ via
@@
where $\xi=3g_{3}(v_{\mathrm{ac}}/v_{\mathrm{zpf}})e^{-i\theta}$ is the three-wave-mixing (3WM) strength, and $K_{\mathrm{eff}}=12(g_{4}-5g_{3}^{2}/\omega_{0})$ is the effective Kerr term (Appendix C). Here $v_{\mathrm{ac}}$ is the ac modulation voltage normalized by the resonator’s zero-point voltage fluctuations $v_{\mathrm{zpf}}$, and $e^{-i\theta}$ arises from the phase of the pump. In Fig. 3 we plot these terms against the external dc bias $v_{0}$.
As is standard in the study of parametric amplifiers [51, 8, 21], Eq. 3, is derived (Appendix C) within the rotating-wave approximation, and, at small enough amplitudes so that higher-order dynamics can be neglected. In our case, however, an expanded expression for $K_{\mathrm{eff}}$, which incorporates second-order corrections from the third-order nonlinearity $g_{3}$, goes to zero at a particular bias point (dashed line in Fig. 3c) [52]. This suggests that there exists a range of external bias $v_{0}$ where the Kerr term can, in principle, be nulled while retaining a significant $\xi$. This cancellation, however, does not reflect an underlying symmetry-breaking structure to our circuit, such as in SNAIL devices [19, 6]. In Appendix D we explore how modifying the circuit geometry impacts the quantities in Eq. 3.
@@
The calculations shown in Fig. 3 indicate that a properly engineered and operated PANDA can be an excellent amplifier with minimal unwanted four-wave-mixing (Kerr) terms that degrades performance. Using the capacitor parameters given in Fig. 1c, and, as an illustrative example, an ac bias of $v_{\mathrm{ac}}=1\penalty 10000\ \,\mathrm{mV}$, we find a maximum three-wave-mixing (3WM) strength of approximately 26 MHz for STO and 9.5 MHz for KTO. Our choice of $v_{\mathrm{ac}}$ translates to pumping with approximately $10^{4}$ photons for both materials; a precise comparison will depend on broader design considerations. Pumping with greater amplitude will increase the 3WM strength, which to the lowest order is proportional to $v_{\mathrm{ac}}$ (Appendix C). Too large a pump field, however, will eventually drive higher-order nonlinearities in a manner dependent on higher-order corrections to Eq. 3, which in general are dependent on dc bias point.
@@
For this design, the optimal 3WM points occur at a dc bias of 9.3 mV for STO and 66 mV for KTO, respectively; modest voltages compatible with superconducting electronics. At these operation points $K_{\mathbf{eff}}$ is of order 0.1 Hz or less for both materials (Fig. 3. For comparison, the Josephson parametric amplifier (JPA) in Ref. [9], operated as a linear amplifier, has $\xi/2\pi\lesssim 25\,\mathrm{MHz}$ and $K_{\mathbf{eff}}/2\pi\approx 50\,\mathrm{kHz}$, and the KIPA in Ref. [21] operates at $\xi/2\pi\lesssim 27\,\mathrm{MHz}$ and $K_{\mathbf{eff}}\approx 0.1\,\mathrm{Hz}$. Our maximum predicted figure-of-merit ratio of $\xi/K_{\mathrm{eff}}\approx 10^{8}$ is much higher than the $\xi/K_{\mathrm{eff}}\lesssim 10^{5}$ ratio typical of Josephson parametric amplifiers (JPA) [53, 6, 8] and is comparable to state-of-the-art KIPAs [20, 21, 30, 24].
@@
Figure 3: (a) Input-output model for a parametric amplifier made from an STO or KTO capacitor. Modulation of the capacitance around $\omega_{p}\approx 2\omega_{0}$ yields degenerate parametric amplification. (b) Magnitude of the three-wave-mixing (3WM) strength $|\xi|/2\pi$ choosing $v_{\mathrm{ac}}=1\,\mathrm{mV}$ (ac bias charge $q_{\mathrm{ac}}=v_{\mathrm{ac}}C(v)$). (c) Magnitude of the effective Kerr strength $K_{\mathrm{eff}}/2\pi$ using a rotating-wave approximation (solid) and one that includes second-order corrections due to $g_{3}$ (dashed). (d,e) Reflection $R(\omega)$ of an STO and KTO device, respectively, plotted for different values of $\xi$ normalized to resonator loss $\kappa/2$. The model (Eq. 4) uses the parameters as given in Fig. 2 and an external quality factor of $Q_{\mathrm{ext}}=100$.
@@
To analyze our circuit as a degenerate parametric amplifier, we consider the reflection of a weak signal tone while the capacitor is charge pumped as previously discussed. Following an input-output formalism applied to $H_{\mathrm{driven}}$ (Eq. 3), the circuit reflection coefficient is [54, 21],
@@
Here $\Delta=\omega_{0}-\omega_{p}/2$ is the detuning of the half-pump frequency from the resonator, and for the remaining analysis we choose $\Delta=0$.
@@
A critical design parameter is the external coupling rate $\kappa_{\mathrm{ext}}$, which should be much larger than the internal loss rate $\kappa_{\mathrm{int}}$ for optimal noise performance. $\kappa_{\mathrm{ext}}$ will be determined by the details of the coupler, whereas $\kappa_{\mathrm{int}}$ depends on the material of choice. Our requirement of introducing both dc and ac bias fields across the capacitor translates to having an inductive coupler. For example, $\kappa_{\mathrm{ext}}$ can be controlled by implementing the coupler as a stepped-impedance filter (see Fig. 7), in further analogy to kinetic inductance amplifiers. Advantageously, this bias scheme avoids challenges with magnetic flux control (including nonuniformity of the bias field and crosstalk between flux lines) that constrain JPAs, e.g. as in Ref. [16]. For the estimates in this paper, we consider an external quality factor $Q_{\mathrm{ext}}=\omega_{0}/\kappa_{\mathrm{ext}}=100$ for both designs, as is typical for superconducting parametric amplifiers [8, 21, 30, 24].
@@
We expect internal loss to be dominated by bulk dielectric loss. The internal quality factor $Q_{\mathrm{int}}=\omega_{0}/\kappa_{\mathrm{int}}=\tan^{-1}{(\delta)}$ is evaluated at $v_{0}$ to be $Q_{\mathrm{int}}\approx 6.1\times 10^{2}$ for STO and $Q_{\mathrm{int}}\approx 7.4\times 10^{3}$ for KTO (from the model plotted in Fig. 4c). These parameters yield $\kappa_{\mathrm{int}}/2\pi=3.4\,\mathrm{MHz}$ and $\kappa_{\mathrm{ext}}/2\pi=20.7\,\mathrm{MHz}$ for the STO design, and $\kappa_{\mathrm{int}}/2\pi=0.7\,\mathrm{MHz}$ and $\kappa_{\mathrm{ext}}/2\pi=48.9\,\mathrm{MHz}$ for the KTO design. (See Fig. 6 for a model of internal loss rate versus resonator frequency.) Both designs are over-coupled ($\kappa_{\mathrm{ext}}/\kappa_{\mathrm{int}}>1$), but the KTO design is more overcoupled due to its lower modeled internal loss. This suggests that PANDAs made from KTO may have better noise performance, as internal loss scales added noise by a factor of $(\kappa_{\mathrm{int}}+\kappa_{\mathrm{ext}})/\kappa_{\mathrm{ext}}$ assuming both the external port and internal bath modes are thermalized to the same temperature [51].